# Controlling Ensemble Variance in Diffusion Models: An Application for Reanalyses Downscaling

### v1

@@ A.1 Properties of element-wise variance
$\mathbb{E}[\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{x}_{t-\Delta t})]$ $\approx\mathbb{E}[\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{m}_{t-\Delta t})+\nabla_{\boldsymbol{m}_{t-\Delta t}}\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{m}_{t-\Delta t})\left(\boldsymbol{x}_{t-\Delta t}-\boldsymbol{m}_{t-\Delta t}\right)]$
$=\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{m}_{t-\Delta t})+\nabla_{\boldsymbol{m}_{t-\Delta t}}\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{m}_{t-\Delta t})\mathbb{E}\left[\left(\boldsymbol{x}_{t-\Delta t}-\boldsymbol{m}_{t-\Delta t}\right)\right]$
$=\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{m}_{t-\Delta t}),$
where in the first equality we used the linearity of the expected value and the observation that $\mathbb{E}[\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{m}_{t-\Delta t})]=\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{m}_{t-\Delta t})$ as $\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{m}_{t-\Delta t})$ is constant with respect to $\boldsymbol{x}_{t-\Delta t}$, while the second equality follows by observing that:
$\mathbb{E}\left[\left(\boldsymbol{x}_{t-\Delta t}-\boldsymbol{m}_{t-\Delta t}\right)\right]=\mathbb{E}[\boldsymbol{x}_{t-\Delta t}]-\mathbb{E}[\boldsymbol{m}_{t-\Delta t}]=\boldsymbol{m}_{t-\Delta t}-\boldsymbol{m}_{t-\Delta t}=\boldsymbol{0}.$
Plugging the equations above into the definition of covariance, we get:
$\boldsymbol{v}(\boldsymbol{x}_{t-\Delta t},\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{x}_{t-\Delta t}))$ $\approx\mathbb{E}\left[(\boldsymbol{x}_{t-\Delta t}-\boldsymbol{m}_{t-\Delta t})(\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{x}_{t-\Delta t})-\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{m}_{t-\Delta t}))^{T}\right]$
$\approx\mathbb{E}\left[(\boldsymbol{x}_{t-\Delta t}-\boldsymbol{m}_{t-\Delta t})(\nabla_{\boldsymbol{m}_{t-\Delta t}}\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{m}_{t-\Delta t})\left(\boldsymbol{x}_{t-\Delta t}-\boldsymbol{m}_{t-\Delta t}\right))^{T}\right]$
$=\nabla_{\boldsymbol{m}_{t-\Delta t}}\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{m}_{t-\Delta t})\mathbb{E}\left[(\boldsymbol{x}_{t-\Delta t}-\boldsymbol{m}_{t-\Delta t})(\left(\boldsymbol{x}_{t-\Delta t}-\boldsymbol{m}_{t-\Delta t}\right))^{T}\right]$
$=\nabla_{\boldsymbol{m}_{t-\Delta t}}\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{m}_{t-\Delta t})\boldsymbol{v}(\boldsymbol{x}_{t-\Delta t}).$
Substituting this result back into the equation for the variance of $\boldsymbol{x}_{t}$, we get:
@@
$\boldsymbol{v}(\boldsymbol{x}_{t})$ $=\frac{\alpha_{t}}{\alpha_{t-\Delta t}}\boldsymbol{v}(\boldsymbol{x}_{t-\Delta t})+c_{t-\Delta t}^{2}\boldsymbol{1}+2\sqrt{\frac{\alpha_{t}}{\alpha_{t-\Delta t}}}c_{t-\Delta t}\nabla_{\boldsymbol{m}_{t-\Delta t}}\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{m}_{t-\Delta t})\boldsymbol{v}(\boldsymbol{x}_{t-\Delta t})$
$=\left(\frac{\alpha_{t}}{\alpha_{t-\Delta t}}\boldsymbol{1}+2\sqrt{\frac{\alpha_{t}}{\alpha_{t-\Delta t}}}c_{t-\Delta t}\nabla_{\boldsymbol{m}_{t-\Delta t}}\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{m}_{t-\Delta t})\right)\boldsymbol{v}(\boldsymbol{x}_{t-\Delta t})+c_{t-\Delta t}^{2}\boldsymbol{1},$
which corresponds to Proposition 3.1 by setting:
@@
$\boldsymbol{F}_{t}:=\frac{\alpha_{t}}{\alpha_{t-\Delta t}}\boldsymbol{1}+2\sqrt{\frac{\alpha_{t}}{\alpha_{t-\Delta t}}}c_{t-\Delta t}\nabla_{\boldsymbol{m}_{t-\Delta t}}\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{m}_{t-\Delta t}),$
$\boldsymbol{g}_{t}:=c_{t-\Delta t}^{2}\boldsymbol{1}.$
A.3 Proof of Theorem 3.2
The proof of the next Theorem is mainly a recursive application of Proposition 3.1 to compute $\boldsymbol{v}_{T}$ from $\boldsymbol{v}_{0}$. Indeed, we recall that:
$v_{t}\approx\boldsymbol{F}_{t}\boldsymbol{v}_{t-\Delta t}+\boldsymbol{g}_{t}.$
Since, as already observed, $\boldsymbol{v}_{0}=\boldsymbol{1}$, then:
$\boldsymbol{v}_{\Delta t}\approx\boldsymbol{F}_{\Delta t}\boldsymbol{1}+\boldsymbol{g}_{\Delta t},$
$\boldsymbol{v}_{2\Delta t}\approx\boldsymbol{F}_{2\Delta t}\boldsymbol{F}_{\Delta t}\boldsymbol{1}+\boldsymbol{F}_{2\Delta t}\boldsymbol{g}_{\Delta t}+\boldsymbol{g}_{2\Delta t},$
$\boldsymbol{v}_{3\Delta t}\approx\boldsymbol{F}_{3\Delta t}\boldsymbol{F}_{2\Delta t}\boldsymbol{F}_{\Delta t}\boldsymbol{1}+\boldsymbol{F}_{3\Delta t}\boldsymbol{F}_{2\Delta t}\boldsymbol{g}_{\Delta t}+\boldsymbol{F}_{3\Delta t}\boldsymbol{g}_{2\Delta t}+\boldsymbol{g}_{3\Delta t},$
$\dots$
@@
$\boldsymbol{v}_{N\Delta t}\approx\left(\prod_{i=1}^{N}\boldsymbol{F}_{i\Delta t}\right)\boldsymbol{1}+\sum_{i=1}^{N}\left(\prod_{k=i+1}^{N+1}\boldsymbol{F_{k\Delta t}}\right)\boldsymbol{g}_{i\Delta t},$
where we defined $\boldsymbol{F}_{(N+1)\Delta t}:=\boldsymbol{I}$ for simplicity, concluding the proof.

### v2

@@ Appendix A Proofs of main results
$\sqrt{\frac{\alpha(t)}{\alpha(t-\Delta t)}}=1-\frac{\alpha^{\prime}(t-\Delta t)}{2\alpha(t-\Delta t)}\,\Delta t+O(\Delta t^{2}),$
and
$c_{t-\Delta t}$ $=\sqrt{1-\alpha(t)}-\sqrt{\alpha(t)}\sqrt{\frac{(1-\alpha(t-\Delta t))}{\alpha(t-\Delta t)}}$
$=\sqrt{1-\alpha(t)}-\sqrt{1-\alpha(t-\Delta t)}+\alpha^{\prime}(t-\Delta t)\frac{\sqrt{1-\alpha(t-\Delta t)}}{2\alpha(t-\Delta t)}\,\Delta t+O(\Delta t^{2})$
$=\frac{\alpha^{\prime}(t-\Delta t)}{2\sqrt{1-\alpha(t-\Delta t)}}\,\Delta t+\alpha^{\prime}(t-\Delta t)\frac{\sqrt{1-\alpha(t-\Delta t)}}{2\alpha(t-\Delta t)}\,\Delta t+O(\Delta t^{2})$
$=\left(\frac{1}{2\sqrt{1-\alpha(t-\Delta t)}}+\frac{\sqrt{1-\alpha(t-\Delta t)}}{2\alpha(t-\Delta t)}\right)\alpha^{\prime}(t-\Delta t)\,\Delta t+O(\Delta t^{2}).$
Plugging these quantities into the definition of $\boldsymbol{F}_{t}$ and $\boldsymbol{g}_{t}$ as in Proposition 3.1 yields:
$\boldsymbol{F}_{t}$ $=\left(\sqrt{\frac{\alpha(t)}{\alpha(t-\Delta t)}}+c_{t-\Delta t}\boldsymbol{J}_{t-\Delta t}\right)^{2}$
$=\Bigg[1-\Bigg(\frac{1}{2\alpha(t-\Delta t)}-\frac{\sqrt{1-\alpha(t-\Delta t)}}{2\alpha(t-\Delta t)}\boldsymbol{J}_{t-\Delta t}+\frac{1}{2\sqrt{1-\alpha(t-\Delta t)}}\boldsymbol{J}_{t-\Delta t}\Bigg)$
$\hskip 220.825pt\cdot\alpha^{\prime}(t-\Delta t)\,\Delta t+O(\Delta t^{2})\Bigg]^{2}$
$=1+\Bigg[-\frac{\alpha^{\prime}(t-\Delta t)}{\alpha(t-\Delta t)}+\boldsymbol{J}_{t-\Delta t}\,\alpha^{\prime}(t-\Delta t)\frac{\sqrt{1-\alpha(t-\Delta t)}}{\alpha(t-\Delta t)}$
@@
$\hskip 180.67499pt+\boldsymbol{J}_{t-\Delta t}\,\frac{\alpha^{\prime}(t-\Delta t)}{\sqrt{1-\alpha(t-\Delta t)}}\Bigg]\Delta t+O(\Delta t^{2})$
@@
$=1+\Bigg(-\partial_{t}\log\alpha(t-\Delta t)+\boldsymbol{J}_{t-\Delta t}\,\alpha^{\prime}(t-\Delta t)\frac{\sqrt{1-\alpha(t-\Delta t)}}{\alpha(t-\Delta t)}$
$\hskip 170.6375pt+\boldsymbol{J}_{t-\Delta t}\,\frac{\alpha^{\prime}(t-\Delta t)}{\sqrt{1-\alpha(t-\Delta t)}}\Bigg)\Delta t+O(\Delta t^{2})$
$=1+\lambda_{t-\Delta t}\Delta t+O(\Delta t^{2}).$
and
$\boldsymbol{g}_{t}$ $=c_{t-\Delta t}^{2}\boldsymbol{v}(\boldsymbol{r}_{t})$
$=\underbrace{\left(\frac{\alpha^{\prime}(t-\Delta t)\sqrt{1-\alpha(t-\Delta t)}}{2\alpha(t-\Delta t)}+\frac{\alpha^{\prime}(t-\Delta t)}{2\sqrt{1-\alpha(t-\Delta t)}}\right)\boldsymbol{v}(\boldsymbol{r}_{t})}_{:=\gamma_{t-\Delta t}}\Delta t+O(\Delta t^{2})$
$=\gamma_{t-\Delta t}\Delta t+O(\Delta t^{2}).$
To show monotonicity of the sequence $\{\boldsymbol{v}_{N\Delta t}\}_{N\in\mathbb{N}}$, note that the above derivation, together with the observation that $\Delta t\to 0$ as $N\to\infty$, implies that:
@@
$\begin{aligned} \log\prod_{i=1}^{N}\boldsymbol{F}_{i\Delta t}&=\sum_{i=1}^{N}\log\left(1+\lambda_{t-\Delta t}\,\Delta t+O(\Delta t^{2})\right)\\[2.84526pt] &\approx\sum_{i=1}^{N}\lambda_{t-\Delta t}\,\Delta t+O(\Delta t^{2})\overset{\Delta t\to 0}{\longrightarrow}\int_{0}^{\infty}\lambda(t)\,dt\\[2.84526pt] \iff\prod_{i=1}^{N}\boldsymbol{F}_{i\Delta t}&\overset{\Delta t\to 0}{\longrightarrow}e^{\int_{0}^{\infty}\lambda(t)\,dt},\end{aligned}$ (20)
where the convergence is monotonic from below due to standard Euler summation arguments. Therefore, by Equation (11), the $N$-th term of the sequence $\{\boldsymbol{v}_{N\Delta t}\}_{N\in\mathbb{N}}$ reads:
$\boldsymbol{v}_{N\Delta t}$ $\approx\left(\prod_{i=1}^{N}\boldsymbol{F}_{i\Delta t}\right)+\sum_{i=1}^{N}\left[\left(\prod_{k=i+1}^{N+1}\boldsymbol{F}_{k\Delta t}\right)\boldsymbol{g}_{i\Delta t}\right],$
and consequently:
$\boldsymbol{v}_{(N+1)\Delta t}$ $\approx\left(\prod_{i=1}^{N+1}\boldsymbol{F}_{i\Delta t}\right)+\sum_{i=1}^{N+1}\left[\left(\prod_{k=i+1}^{N+2}\boldsymbol{F}_{k\Delta t}\right)\boldsymbol{g}_{i\Delta t}\right]$
[…]
$=\left(\prod_{i=1}^{N}\boldsymbol{F}_{i\Delta t}\right)+\sum_{i=1}^{N}\left[\left(\prod_{k=i+1}^{N+1}\boldsymbol{F}_{k\Delta t}\right)\boldsymbol{g}_{i\Delta t}\right]+\boldsymbol{g}_{(N+1)\Delta t}$
$=\boldsymbol{v}_{N\Delta t}+\boldsymbol{g}_{(N+1)\Delta t}$
$\geq\boldsymbol{v}_{N\Delta t}\quad\text{(since }\boldsymbol{g}_{(N+1)\Delta t}\geq 0\text{)},$
proving monotonicity. Arguments similar to (20) also show that:
$\boldsymbol{v}_{N\Delta t}\overset{\Delta t\to 0}{\longrightarrow}e^{-\int_{0}^{T}\lambda(s)\,ds}+\int_{0}^{T}e^{-\int_{0}^{s}\lambda(\kappa)\,d\kappa}\gamma(s)\,ds,$
concluding the proof. ∎
Appendix B Extension to the $n$-Dimensional Case
In the main text, we presented the complete theoretical analysis for the scalar case ($n=1$), which already provides an intuitive understanding of the evolution of the variance across DDIM timesteps. For completeness, in this Appendix we report the extended derivation for the general $n$-dimensional case, where $\boldsymbol{x}_{t}\in\mathbb{R}^{n}$. The same reasoning applies, but with the variance replaced by the full covariance matrix, and the Jacobian $\boldsymbol{J}_{t}$ promoted to a matrix. We then show that the total variance, defined as the trace of the covariance matrix, follows an analogous affine recursion and converges monotonically to an exponential-affine limit.
Proposition B.1.
Let $\boldsymbol{m}_{t}:=\mathbb{E}[\boldsymbol{x}_{t}]$ and $\boldsymbol{\Sigma}_{t}:=Var(\boldsymbol{x}_{t})$. Then, under the first-order linearization of the DDIM update rule, it holds:
$\boldsymbol{\Sigma}_{t}\approx\boldsymbol{A}_{t}\boldsymbol{\Sigma}_{t-\Delta t}\boldsymbol{A}_{t}^{T}+c_{t-\Delta t}^{2}Var(\boldsymbol{r}_{t}),$ (21)
where:
$\boldsymbol{A}_{t}:=\sqrt{\frac{\alpha_{t}}{\alpha_{t-\Delta t}}}\,\boldsymbol{I}+c_{t-\Delta t}\boldsymbol{J}_{t-\Delta t},$
and $\boldsymbol{r}_{t}$ is a random vector independent of $\boldsymbol{x}_{t-\Delta t}$.
Proof.
Starting from the DDIM step
$\boldsymbol{x}_{t}=\sqrt{\frac{\alpha_{t}}{\alpha_{t-\Delta t}}}\boldsymbol{x}_{t-\Delta t}+c_{t-\Delta t}\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{x}_{t-\Delta t}),$
we linearize $\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{x}_{t-\Delta t})$ around $\boldsymbol{m}_{t-\Delta t}:=\mathbb{E}[\boldsymbol{x}_{t-\Delta t}]$:
$\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{x}_{t-\Delta t})\approx\boldsymbol{E}_{t-\Delta t}+\boldsymbol{J}_{t-\Delta t}(\boldsymbol{x}_{t-\Delta t}-\boldsymbol{m}_{t-\Delta t})+\boldsymbol{r}_{t},$
where $\boldsymbol{E}_{t-\Delta t}=\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{m}_{t-\Delta t})$, $\boldsymbol{J}_{t-\Delta t}$ is the Jacobian of $\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}$ evaluated at $\boldsymbol{m}_{t-\Delta t}$, and $\boldsymbol{r}_{t}$ is the Taylor remainder, independent of $\boldsymbol{x}_{t-\Delta t}$. Plugging this into the DDIM step yields:
$\boldsymbol{x}_{t}$ $\approx\left(\sqrt{\frac{\alpha_{t}}{\alpha_{t-\Delta t}}}\boldsymbol{I}+c_{t-\Delta t}\boldsymbol{J}_{t-\Delta t}\right)\boldsymbol{x}_{t-\Delta t}+c_{t-\Delta t}\boldsymbol{r}_{t}+c_{t-\Delta t}\boldsymbol{E}_{t-\Delta t}-c_{t-\Delta t}\boldsymbol{J}_{t-\Delta t}\boldsymbol{m}_{t-\Delta t}.$
Since the last two terms are constant with respect to $\boldsymbol{x}_{t-\Delta t}$, their variance is zero. Denoting $\boldsymbol{A}_{t}:=\sqrt{\frac{\alpha_{t}}{\alpha_{t-\Delta t}}}\boldsymbol{I}+c_{t-\Delta t}\boldsymbol{J}_{t-\Delta t}$, taking the variance on both sides gives:
$Var(\boldsymbol{x}_{t})$ $\approx Var\big(\boldsymbol{A}_{t}\boldsymbol{x}_{t-\Delta t}+c_{t-\Delta t}\boldsymbol{r}_{t}\big)$
$=\boldsymbol{A}_{t}Var(\boldsymbol{x}_{t-\Delta t})\boldsymbol{A}_{t}^{T}+c_{t-\Delta t}^{2}Var(\boldsymbol{r}_{t}),$
which is exactly (21). ∎
To obtain a scalar measure of the overall uncertainty, we now consider the total variance $\boldsymbol{v}(\boldsymbol{x}_{t}):=\operatorname{Tr}(\boldsymbol{\Sigma}_{t})$. Using $\operatorname{Tr}(\boldsymbol{A}\boldsymbol{B})=\operatorname{Tr}(\boldsymbol{B}\boldsymbol{A})$ and $\operatorname{Tr}(\boldsymbol{A}^{T}\boldsymbol{A})=\|\boldsymbol{A}\|_{F}^{2}$, we can derive the following corollary.
Corollary B.2.
The total variance $\boldsymbol{v}_{t}:=\boldsymbol{v}(\boldsymbol{x}_{t})$ approximately satisfies the scalar affine recursion:
$\boldsymbol{v}_{t}\approx F_{t}\,\boldsymbol{v}_{t-\Delta t}+G_{t},$ (22)
where:
$F_{t}:=\frac{1}{n}\operatorname{Tr}(\boldsymbol{A}_{t}^{T}\boldsymbol{A}_{t})=\frac{1}{n}\|\boldsymbol{A}_{t}\|_{F}^{2}>0,$
$G_{t}:=c_{t-\Delta t}^{2}\,\boldsymbol{v}(\boldsymbol{r}_{t})\geq 0.$
Equation (22) generalizes the scalar recurrence (11) to arbitrary dimension $n$, preserving the same structure. By iteratively applying this recursion, we obtain a closed-form expression for $\boldsymbol{v}_{T}$, analogous to Theorem 3.2 in the main text, and a continuous-time limit with exponential convergence properties.
Theorem B.3.
Let $\boldsymbol{v}_{t}$ satisfy (22) and assume that $\alpha(t)$, $\boldsymbol{J}_{t}$, and $\boldsymbol{r}_{t}$ are continuously differentiable in $t$. Then, as $\Delta t\to 0$, it holds:
$\lim_{\Delta t\to 0}\boldsymbol{v}_{T}=e^{-\int_{0}^{T}\lambda(s)ds}+\int_{0}^{T}e^{-\int_{0}^{s}\lambda(\kappa)d\kappa}\,\gamma(s)\,ds,$ (23)
where:
$\lambda(t)=-\partial_{t}\log\alpha(t)+\frac{1}{n}\operatorname{Tr}(\boldsymbol{J}_{t})\,\alpha^{\prime}(t)\left(\frac{\sqrt{1-\alpha(t)}}{\alpha(t)}+\frac{1}{\sqrt{1-\alpha(t)}}\right),$
$\gamma(t)=\left(\frac{\alpha^{\prime}(t)\sqrt{1-\alpha(t)}}{2\alpha(t)}+\frac{\alpha^{\prime}(t)}{2\sqrt{1-\alpha(t)}}\right)\boldsymbol{v}(\boldsymbol{r}_{t}).$
Proof.
The proof follows exactly the same steps as Theorem 3.2 for the scalar case. Taylor expanding $\sqrt{\alpha(t)}$ and $\sqrt{1-\alpha(t)}$ around $t-\Delta t$, and using:
$\sqrt{\frac{\alpha(t)}{\alpha(t-\Delta t)}}$ $=1-\frac{\alpha^{\prime}(t-\Delta t)}{2\alpha(t-\Delta t)}\Delta t+O(\Delta t^{2}),$
$c_{t-\Delta t}$ $=\left(\frac{1}{2\sqrt{1-\alpha(t-\Delta t)}}+\frac{\sqrt{1-\alpha(t-\Delta t)}}{2\alpha(t-\Delta t)}\right)\alpha^{\prime}(t-\Delta t)\Delta t+O(\Delta t^{2}),$
yields:
$\boldsymbol{A}_{t}$ $=\boldsymbol{I}+\Delta t\left[-\frac{1}{2}\partial_{t}\!\log\alpha(t-\Delta t)\,\boldsymbol{I}+q_{t-\Delta t}\,\boldsymbol{J}_{t-\Delta t}\right]+O(\Delta t^{2}),$
where $q_{t-\Delta t}=\frac{\alpha^{\prime}(t-\Delta t)}{2\alpha(t-\Delta t)}\sqrt{1-\alpha(t-\Delta t)}+\frac{\alpha^{\prime}(t-\Delta t)}{2\sqrt{1-\alpha(t-\Delta t)}}.$ Therefore:
$F_{t}$ $=\frac{1}{n}\operatorname{Tr}(\boldsymbol{A}_{t}^{T}\boldsymbol{A}_{t})=1+\lambda(t-\Delta t)\Delta t+O(\Delta t^{2}),$
$G_{t}$ $=c_{t-\Delta t}^{2}\,\boldsymbol{v}(\boldsymbol{r}_{t})=\gamma(t-\Delta t)\Delta t+O(\Delta t^{2}),$
with $\lambda$ and $\gamma$ defined as above. Unrolling the recursion (22) leads to a product–sum expression identical to (13), and standard Euler-sum arguments show that the discrete products $\prod F_{t}$ converge monotonically from below to $e^{\int\lambda}$, while the additive term converges to the integral in (23). Monotonicity of $\{\boldsymbol{v}_{N\Delta t}\}_{N\in\mathbb{N}}$ follows from the same argument used in the scalar case: since $F_{t}>0$ and $G_{t}\geq 0$, each refinement step increases $\boldsymbol{v}_{t}$, completing the proof. ∎